OpenStudy (anonymous):

How do you find the x intercepts in a quadratic equation? I know how to find the vertex but my teacher didn't explain it very well on finding the x intercepts.

OpenStudy (anonymous):

set the quadric equal to zero and solve for \(x\) there are many method for solving 1) factor if you can 2) complete the square, easiest if leading coefficient is 1 and middle term is even 3) quadratic formula, it always works \[ax^2+bx+c=0\implies x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]